A microwave oven cooks food with microwaves of wavelength 14.9 cm. What is the frequency of these microwaves?
Answer #1

We know that c= f\lambda

We have c = speed of light= 3 ×10⁸ m/sec

\lambda = 11.8cm = 11.8 × 10-2 m

Thus f = c/\lambda

f = 3×10⁸ /11.8×10-2 s-1

f = 2.54 × 10⁹Hz


answered by: ANURANJAN SARSAM